OpenStudy (anonymous):

Some questions , I'll post in a second , please help:

OpenStudy (anonymous):

Q1/ quantity of charge in coulombs passing though surface of area 2 cm^2 varies with time according to the equation q=4t^3+5t+6 where t is in seconds a)what is the instantaneous current through the surface at t=1s b)what is the value of current density a) I did not know the answer :( b) I think we should use j=I/A but I is missing I should get it from a) :(

OpenStudy (kc_kennylau):

Can't you even write the topic out instead of "help"? P.S. that's just the same.

OpenStudy (anonymous):

sorry :(

OpenStudy (kc_kennylau):

a) just plug in t=1 :)

OpenStudy (anonymous):

Q2/ A 0.9 v potential difference maintained across a 1.5 m length of tungsten wire that has a cross sectional area of 0.6mm^2 what is the current in the wire ? I think we should use I=V/R , but from where can I get R ?

OpenStudy (kc_kennylau):

Do you have the resistance of tungsten wire?

OpenStudy (anonymous):

Can i get r from r/l = resistance of the element / 0.6*10^-6 ?

OpenStudy (anonymous):

I can look it up , it doesn't matter but is my solution right ?

OpenStudy (kc_kennylau):

where did you get 0.6*10^-6?

OpenStudy (anonymous):

Sorry * 10^-3

OpenStudy (kc_kennylau):

where did you get it?

OpenStudy (anonymous):

Q4/ 10V battery is connected to a 120 ohmm resistor, neglecting the internal resistance of the battery calculate the power delivred to the resistor ? Should I find the current first , =v/r and then use p=i^2 * r or p=i *v ?

OpenStudy (anonymous):

from mm^2 to m^2

OpenStudy (kc_kennylau):

Q4 doesn't have a question lol read it again

OpenStudy (anonymous):

:D it says calculate the power that is delivered to the resistor

OpenStudy (kc_kennylau):

What are you given? (this is the guiding question lol)

OpenStudy (anonymous):

Hey gave delta V also gave me the value of the resistor

OpenStudy (anonymous):

V = 10v ,, R = 120 ohmm

OpenStudy (kc_kennylau):

So it's voltage and resistance? Use P=V^2/R :D

OpenStudy (kc_kennylau):

https://en.wikipedia.org/wiki/Electric_power

OpenStudy (anonymous):

can I find the current first by I = V/R and then use P = I ^2 * R ?

OpenStudy (kc_kennylau):

ofc :)

OpenStudy (kc_kennylau):

In fact P=I^2R is derived from P=V^2/R :D

OpenStudy (kc_kennylau):

https://en.wikipedia.org/wiki/Electric_power

OpenStudy (anonymous):

Ah man I hate this subject :( good thing it's the last time I'll study this subject :D Thank you so much .

OpenStudy (kc_kennylau):

no problem :) good luck with your study onwards :D